OpenStudy (anonymous):

1:Write the equation in slope intercept form. What are the slope and y-intercept? -9x+10y=-9

OpenStudy (anonymous):

@tyler_peters here yo go! :)

OpenStudy (anonymous):

Simplify -9x + 10y = -9 Solve -9x + 10y = -9 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-10y' to each side of the equation. -9x + 10y + -10y = -9 + -10y Combine like terms: 10y + -10y = 0 -9x + 0 = -9 + -10y -9x = -9 + -10y Divide each side by '-9'. x = 1 + 1.111111111y Simplify x = 1 + 1.111111111y

OpenStudy (anonymous):

Too late I already got the question wrong but thank you for answering.